You have 26 red cards and 26 black cards. You draw 2 cards and draw another 2 cards without replacing the cards. What is the probability you will draw red cards?
Answer by Boreal(15235)

It isn't clear how many red cards are desired. If all 4 are red, the first has a probability of 1/2, the second 25/51, the third 24/50 (12/25) and the fourth 23/49. That probability is 13800/249900=0.055
This is intuitively correct since if independent it would be (1/2)^4=1/16=0.0625, which is a little higher, since each probability would be now be 1/2, rather than a little less.
If the question has a different interpretation, please re-submit.